Light of wavelength 587.5 nm illuminates a slit of width 0.75 mm.

(a) At what distance from the slit should a screen be placed if the first minimum in the diffraction pattern is to be 0.85mm from the central maximum?

(b) Calculate the width of the central maximum.


1
Expert's answer
2021-02-25T11:38:15-0500

Answer

a)

"\lambda=587.5\times10^{-9}m"

a="0.75\times10^{-9}m"

"y_1=0.85\times10^{-3}m"

distance from the slit should a screen be placed if the first minimum in the diffraction pattern is to be 0.85mm from the central maximum

"L=y_1\frac{a}{\lambda}"

="\frac{0.85\times10^{-3}\times0.75\times10^{-9}}{587.5\times10^{-9}}"

=1.1m


b) the width of the central maximum.

"X=\frac{2D\lambda}{d}"

"=\frac{2\times1.1\times587.5\times10^{-9}}{0.85\times10^{-3}}\\=1.5mm"